A type of milk contains 8% water.What quantity of pure milk should be added to 20litres of this milk to reduce water to 5%?
Answer by robertb(5830)

Let x = # of liters of pure milk to be added.
==> 0.08*20 + 0.00x = 0.05(20 + x)
<==> 1.6 = 1 + 0.05x
==> 0.6 = 0.05x
==> x = 12 liters of pure milk should be added.